A room is 12m long and 8m broad.Find the cost of carpeting it at the rate of Rs.6.5/m when the carpet is 75cm broad
Answer by glabow(165)

We can start by assuming that the carpet can be installed running either way. (If not, the answer will be different.) Since the carpetting is .75 meters wide, there will have to be 12m / .75m strips of carpet to cover the floor. Each strip will be 8m long. So the entire length of the carpeting is
12m / .75 m * 8 m = 128m
At 6.5 per meter it will cost 128m * 6.5 /m = 832.
If the carpet must be laid with its width running with the width of the room, then we need 8m / .75 m strips of carpet to cover the floor. 8m /.75 m = 10.666... strips, each 12 m long. But we have to buy 11 strips and throw some away. 11 strips * 12 m/strip = 132 m. So the cost in this case is 132m * 6.5 /m = 858.
Take your pick.